OpenStudy (anonymous):

Find the derivative of f(x) = -12x2 + 9x at x = 6.

OpenStudy (anonymous):

differentiate it using this \[\frac{ d }{ dx }(cx^n) = cnx ^{n - 1}\]

OpenStudy (anonymous):

i thought the formula was \[f(x + h) - f(x) / h\]

OpenStudy (anonymous):

Thats for the limit isn't it?\[\lim_{h \rightarrow 0}\frac{ f(x + h) - f(x) }{ h }\]

OpenStudy (anonymous):

oh yeah. so i use the other formula for the derivative?

OpenStudy (anonymous):

Use the previous formula

OpenStudy (anonymous):

im still confused, im not sure where to plug the numbers into

OpenStudy (anonymous):

\[f(x) = -12x^2 + 9x\]\[f'(x) = -12(2)x + 9\]\[f'(x) = -24x + 9\]Now substitute x = 6 into the derivative.

OpenStudy (anonymous):

153

OpenStudy (anonymous):

???

OpenStudy (anonymous):

24(6) + 9 ?

OpenStudy (anonymous):

it's -24....

OpenStudy (anonymous):

-135